OpenStudy (anonymous):

what is -0.02 as a fraction

OpenStudy (danjs):

what place is the 2 in ?

OpenStudy (anonymous):

ones?

OpenStudy (danjs):

that is the hundreths place. a easy way to make a fraction is to put that many over 100. -2/100

OpenStudy (anonymous):

oh well my actual question was 9/8n= -54

OpenStudy (danjs):

multiply both sides of that by 8/9 8/9 * 9/8 = 1

OpenStudy (anonymous):

ok

OpenStudy (danjs):

that will give you n = -54 * (8/9)

OpenStudy (anonymous):

umm ok

OpenStudy (anonymous):

im not sure

OpenStudy (anonymous):

@DanJS

OpenStudy (anonymous):

48?

OpenStudy (anonymous):

n=48

OpenStudy (danjs):

it will be a negative number, you have -54 times 8/9, negative times a positive = a negative

OpenStudy (danjs):

\[\frac{ 9 }{ 8 }n = -54\] \[\frac{ 8 }{ 9 }*\frac{ 9 }{ 8 }n = -54\frac{ 8 }{ 9 }\]\[1n = \frac{ -54*8 }{ 9 }\]

OpenStudy (anonymous):

-48